# DiMo-GUI: Advancing Test-time Scaling in GUI Grounding via Modality-Aware Visual Reasoning

### v1

@@ Abstract
The corresponding author.The work was done during the first author’s internship at vivo Company.
@@ 2.1 GUI Agents
Recent years have witnessed significant advances in GUI automation driven by large language models (LLMs). Early GUI agents predominantly focused on web interactions Nakano et al. (2022); Hong et al. (2024) and have gradually expanded to mobile Zhang et al. (2023); Wang et al. (2024a) and desktop environments Zhang et al. (2024). A fundamental challenge across these applications is precise element localization. Traditional approaches relied on structured information like XML and DOM trees Zhang et al. (2023), but faced limitations in accessibility and information redundancy. Alternative methods using OCR Du et al. (2020) or detection models Liu et al. (2024a) introduced additional computational overhead. Recent advances in multimodal large language models (MLLMs) have enabled direct GUI element localization Hong et al. (2024); Cheng et al. (2024); Lin et al. (2024), partially bridging the visual perception gap. Tang et al. (2025) introduces a dual-system framework that combines fast prediction with systematic analysis to provide robust GUI foundation. OS-Atlas Wu et al. (2024) and UGround Gou et al. (2024) created large datasets and trained models to handle out-of-distribution tasks.  Zhou et al. (2025); Lee et al. (2025a); Yuan et al. (2025); Xia and Luo (2025) explored improving grounding performance using reinforcement learning.
@@ 2.2 Test-time scaling
Test-time scaling dynamically adjusts computational resources during inference to enhance model performance, with recent studies showing it can outperform increased train-time computation through strategies like best-of-N sampling and external verification Snell et al. (2024); Lee et al. (2025b); Hosseini et al. (2024). In localization tasks, test-time scaling has also been framed as a search problem Wu and Xie (2024). Inspired by its success in LLMs, similar techniques have been applied to GUI agents, such as leveraging action histories Zhang and Zhang (2023), gathering external information Nakano et al. (2022), zooming in and searching Nguyen (2024), and adaptively refining focus regions Luo et al. (2025).
@@
Input: Full-resolution GUI image $I$, instruction $Q$
@@
[…]
@@
[…]
@@ 9 $C^{*}\leftarrow\texttt{Select}(C_{\text{text}},C_{\text{icon}},I,Q)$
11 Function: DynamicGrounding($I$, $Q$, modality $m$)
@@ 3.1 Dynamic Grounding Mechanism
The number of iterations in the zooming process plays a critical role in determining the final grounding performance. Since different GUI screenshots and user instructions vary in complexity, it is evident that a fixed number of iterations is not optimal for all cases. To address this, we introduce a dynamic iteration mechanism that allows the model to autonomously decide whether to DiMo-GUI early during the progressive narrowing process. This approach not only reduces unnecessary iterations and improves inference efficiency but also prevents the model from "overthinking"—i.e., drifting into incorrect regions after having already located the correct target. Specifically, the method determines whether to continue zooming based on the spatial distance between the inference results before and after zooming. If the spatial distance between the predicted coordinates is smaller than one-sixth of the diagonal length of the pre-zoom image, it indicates that the target region has been localized with sufficient precision. In this case, further zooming is DiMo-GUIped, and the final coordinates are returned as the result. The above process is described as DiMo-GUICondition($C_{t}$, $t$) in the algorithm 1, which decides whether to DiMo-GUI dynamic zooming in the t iteration based on the predicted coordinates $C_{t}$. Additionally, to prevent excessive zooming, we set an upper limit max_iters of seven zooming iterations.

### v2

@@ Abstract
The corresponding author.
@@ 2.1 GUI Agents
Recent years have witnessed significant advances in GUI automation driven by large language models (LLMs). Early GUI agents predominantly focused on web interactions Nakano et al. (2022); Hong et al. (2024) and have gradually expanded to mobile Zhang et al. (2023); Wang et al. (2024a) and desktop environments Zhang et al. (2024). A fundamental challenge across these applications is precise element localization. Traditional approaches relied on structured information like XML and DOM trees Zhang et al. (2023), but faced limitations in accessibility and information redundancy. Alternative methods using OCR Du et al. (2020) or detection models Liu et al. (2024a) introduced additional computational overhead. Recent advances in multimodal large language models (MLLMs) have enabled direct GUI element localization Hong et al. (2024); Cheng et al. (2024); Lin et al. (2024), partially bridging the visual perception gap. Tang et al. (2025) introduces a dual-system framework that combines fast prediction with systematic analysis to provide robust GUI foundation. OS-Atlas Wu et al. (2024) and UGround Gou et al. (2024) created large datasets and trained models to handle out-of-distribution tasks.  Zhou et al. (2025); Lee et al. (2025a); Yuan et al. (2025); Xia and Luo (2025) explored improving grounding performance using reinforcement learning.  Tao et al. (2025) proposes a framework and method to diagnose and reduce localization errors in MLLMs for GUI interaction, improving interpretability and robustness.
@@ 2.2 Test-time scaling
Test-time scaling dynamically adjusts computational resources during inference to enhance model performance, with recent studies showing it can outperform increased train-time computation through strategies like best-of-N sampling and external verification Snell et al. (2024); Lee et al. (2025b); Hosseini et al. (2024). In localization tasks, test-time scaling has also been framed as a search problem Wu and Xie (2024). Inspired by its success in LLMs, similar techniques have been applied to GUI agents, such as leveraging action histories Zhang and Zhang (2023), gathering external information Nakano et al. (2022), zooming in and searching Nguyen (2024), and adaptively refining focus regions Luo et al. (2025).  Ge et al. (2025) proposes Multi-Region Fusion Decoding (MRFD), a training-free method that reduces hallucinations in LVLMs by leveraging inter-region consistency to improve factual grounding.
@@
Input: Full-resolution GUI image II, instruction $Q$
@@
3 $C_{\text{text}}\leftarrow\texttt{DynamicGrounding}(I,Q,\text{``text''})$
@@
6 $C_{\text{icon}}\leftarrow\texttt{DynamicGrounding}(I,Q,\text{``icon''})$
@@ 9 $C^{*}\leftarrow\texttt{Select}(C_{\text{text}},C_{\text{icon}},I,Q)$
11 Function: DynamicGrounding(II, $Q$, modality $m$)
@@ 3.1 Dynamic Grounding Mechanism
The number of iterations in the zooming process plays a critical role in determining the final grounding performance. Since different GUI screenshots and user instructions vary in complexity, it is evident that a fixed number of iterations is not optimal for all cases. To address this, we introduce a dynamic iteration mechanism that allows the model to autonomously decide whether to DiMo-GUI early during the progressive narrowing process. This approach not only reduces unnecessary iterations and improves inference efficiency but also prevents the model from "overthinking"—i.e., drifting into incorrect regions after having already located the correct target. Specifically, the method determines whether to continue zooming based on the spatial distance between the inference results before and after zooming. If the spatial distance between the predicted coordinates is smaller than one-sixth of the diagonal length of the pre-zoom image, it indicates that the target region has been localized with sufficient precision. In this case, further zooming is stopped, and the final coordinates are returned as the result. The above process is described as StopCondition($C_{t}$, $t$) in the algorithm 1, which decides whether to stop dynamic zooming in the t iteration based on the predicted coordinates $C_{t}$. Additionally, to prevent excessive zooming, we set an upper limit max_iters of seven zooming iterations.
@@ 6 Limitations
7 Acknowledgments
The work is partially supported by the NSF of the United States Grant CRII 2451683, an NVIDIA Academic Grants Program, University of California at Merced, and a UC Merced Faculty Research Award. The views and conclusions are those of the authors and should not reflect the official policy or position of the U.S. Government.